OpenStudy (anonymous):

WILL MEDAL Solve each equation. Check your answer: Z - 4 2/3= 2 2/3 THANKS SO MUCH

OpenStudy (anonymous):

By 4 2/3, you mean 4 and 2/3 as a mixed number?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

First, add 4 2/3 to each side like this z - 4 2/3 = 2 2/3 + 4 2/3 + 4 2/3 Z = (answer here)

OpenStudy (anonymous):

What do you think the answer would be?

OpenStudy (anonymous):

just a second

OpenStudy (anonymous):

i dont get the right answer. please help some more. i tried. am i doing something wrong. this is what i did on my texas calculator: 4(2/3) + 4(2/3) and i also did 2(2/3) + 4(2/3). is that right?

OpenStudy (anonymous):

Both ways you did that would be correct. Give me a second and let me work out how to do it so I can explain it better :)

OpenStudy (anonymous):

1) Convert the mixed number to an improper fraction 2*3=6+2=8 // 4*3=12+2=14 \[\frac{ 8 }{ 3 } + \frac{ 14 }{ 3 } = \frac{ 22 }{ 3 }\]

OpenStudy (anonymous):

22 ÷ 3 = 7 1/3

OpenStudy (anonymous):

that gives me -1/3 though as my answer if i do 7(1/3) - 4(2/3)

OpenStudy (anonymous):

is that what you wanted me to do

OpenStudy (anonymous):

How did you get -1/3?

OpenStudy (anonymous):

i did 7(1/3) - 4(2/3)

OpenStudy (anonymous):

Why did you subtract 4(2/3) from 7(1/3)? I'm not being harsh, just curious :)

OpenStudy (anonymous):

i thought that 7(1/3) was what Z equaled because of what you said. i must have read your answer wrong.

OpenStudy (anonymous):

Z would just be: Z = 7(1/3), no subtracting or anything. Do you understand what I did, though?

OpenStudy (anonymous):

yes. but if you look at my first question you will see that i have to subtract to check the answer.

OpenStudy (anonymous):

Ahh, sorry. I did not see that. That is correct, sorry for getting so confused xD

OpenStudy (anonymous):

so is your answer right?

OpenStudy (anonymous):

Yes it is.

OpenStudy (anonymous):

thank you so helpful